hello, i have a quiz coming up and i need help :) please SHOW work and EXPLAIN how to do the problem :
-3/5 + -3/4 - 7/8

Answer by rfer(16322)

find common denominator first
-24/40-30/40-35/40=-89/40= -2 9/40
The + - 3/4 becomes a minus
minus times a plus is a minus
reduce to lowest form -89/40